# On Developers’ Self-Declaration of AI-Generated Code: An Analysis of Practices

### v2

@@ 1. Introduction
Despite their usefulness and high potential (Wang et al., 2024; Liang et al., 2024), previous studies have shown that there are many quality (Yetiştiren et al., 2023; Liu et al., 2024b), correctness (Tambon et al., 2025; Vatai et al., 2024; Liu et al., 2024c), security (Pearce et al., 2022; Majdinasab et al., 2024; Fu et al., 2025) and ethical  (Stalnaker et al., 2024) challenges associated with AI-generated code (i.e., code that is primarily generated with AI coding assistance tools). These challenges stress the need to distinguish AI-generated code from other human-written code (Xu et al., 2025; Idialu et al., 2024) for code review and quality assurance purposes (Xu et al., 2025). To this end, we seek to study whether developers self-declare AI-generated code, explore the characteristics of self-declaration comments, and explore how and why developers self-declare (or do not) AI-generated code. We derive our definition of self-declared AI-generated code from the well-established definition of self-admitted technical debt (Potdar and Shihab, 2014): a self-declared AI-generated code is code that has been explicitly acknowledged in the code comments by the developer as automatically generated by AI tools.
@@ 3. Research Methodology
Our study employs a mixed-methods approach comprising two main phases: (1) an analysis of self-declaration comments from AI-generated code collected from open-sourced projects, and (2) a practitioners’ survey with practitioners regarding the practices of self-declaring AI-generated code. We first collected data about AI-generated code from real-world projects by mining GitHub repositories to answer our research question of how (RQ1) developers self-declare AI-generated code. We followed that with a practitioner survey to validate the results from the mining study (RQ1) while addressing the research question of why (RQ2) developers self-declare AI-generated code. This methodological triangulation strengthens the validity of our findings by complementing the mining study results with developers’ perspectives from the practitioners’ survey. Figure 2 shows an overview of our research methodology.
@@ 1 Tool Names Only ChatGPT, GitHub Copilot, Code Llama, Amazon CodeWhisperer
In the pilot search phase, we collected a significant number of self-declared AI-generated code using search terms that included the names of the AI code generation tools and phrases specific to code generation. Therefore, out of the total 23 search terms finally selected for our data collection, 20 were based on Strategy 3 (see Table 1). The list included only 2 “thanking” terms from Strategy 4 because we also found relevant results using these two terms during the pilot phase. We only included one general search term without the AI tool name (Strategy 2) because we found limited results for such search terms. We could not find any results using the terms from Strategy 1, and we did not include any search terms from this strategy. Based on the results of the pilot search, 22 search terms included the names of the AI code generation tools. 17 out of the 22 search terms included names of ChatGPT and GitHub Copilot, and only 5 search terms included names of other tools.. The list of search terms is provided in Table 2, which also shows the number of AI-generated code snippets included in our dataset.
@@ 3.2.2. Search and Filtration Process
We initially curated slightly over 1,200 code files from public GitHub repositories, and after applying our inclusion and exclusion criteria (see Table 3), we retained 613 code files from 586 unique repositories for data analysis. To ensure a diverse and non-redundant dataset for data analysis, we included a few representative files from repositories containing multiple files with AI-generated code. This was necessary as several repositories contained many files with highly similar self-declaration comments. For instance, the repository KV6002-NUMUNCHIES alone contained 65 files with similar self-declaration comments with small variations. On the contrary, for the analysis of the distribution of AI-generated code in RQ1.3, we recorded the total number of files containing self-declared AI-generated code snippets in “Total files with SDC” column (code-details.xlsx) in our dataset (Kashif et al., 2025).
@@ 3.2.4. Data Analysis
Initial Coding. We performed the initial coding by identifying and labeling essential ideas and concepts from self-declaration comments. The first author carefully read each self-declaration comment to extract codes. We iteratively refined the code as we examined more data. The other co-authors continuously reviewed the codes and shared their feedback to improve the reliability of the coding process. The analysis was done using MAXQDA [^2] https://www.maxqda.com/ to organize and manage the coding results.
@@ 3.3.3. Data Analysis
Analyzing Responses. We analyzed all responses to the survey questionnaire in Table 5 using a combination of quantitative and qualitative analysis methods. For the first nine closed-ended questions, we employed descriptive statistics to ascertain the frequency of various responses, which provides an overview of the participants’ demographics and their usage of AI-generated code. For the remaining three open-ended questions (SQ10, SQ11, SQ12), we employed thematic analysis and utilized the inductive approach for analysis (Braun and Clarke, 2006). Specifically, we first performed initial coding and then constantly compared the codes to develop high-level themes. The first author performed the initial coding by examining each valid response to extract and summarize the core ideas articulated in participants’ responses regarding their reasons to self-declare AI-generated code.
@@ 3.4. Mining Study Overview
Commit history. Figure 6 shows the commit history of the repositories that contain self-declared AI-generated code. We categorized repositories into six commit count ranges: from 1 – 50 to $\geq$ 5000. In total, we collected data from 586 repositories. 194 (33.1%) repositories are in the 100 - 500 commits range, indicating reasonable activity in these projects. Moreover, 137 repositories (23.4%) contain commit counts ranging from 1 to 50, showing certain inactivity in these projects. The projects containing more than 500 commits account for 136 (23.2%) repositories.
@@ 4.1. RQ1: How do developers self-declare AI-generated code in their projects?
1) Simple self-declaration. We identified 214 (34.9%) self-declaration comments as simple self-declaration comments. These comments self-declare AI-generated code in a few words, simply acknowledging the code is AI-generated. In most cases (over 96%), the comment will also include the name of the code generation tool used. The example (Comment 1) in Table 7 shows a developer mentioning that Copilot was used to generate the code. Only 3.3% of these comments just self-declared the AI-generated code without mentioning the tool name. For example, we can see in Comment 3 in Table 7 that a developer simply self-declared a function as AI-generated without mentioning any further details. In some self-declaration comments (4%) also acknowledged the simultaneous use of multiple AI code generation tools. In Comment 2 in Table 7, a developer mentioned GitHub Copilot and ChatGPT simultaneously for code generation.
@@ 7. Written by [developer name], 4/9/2024 with the assistance of GitHub Copilot (comment)
Trend of self-declaring AI-generated code among developers. We visualized the dates when self-declaration comments were added to the files, covering the period from January 2021 to December 2024. We observed an increasing trend in self-declaring AI-generated code among developers, as shown in Figure 14. The trend exhibits continuous growth, punctuated by sudden spikes. A noticeable increase is observed starting in November 2022, coinciding with the introduction of the GPT-3.5 model. Similarly, we observed another significant rise in March 2023, when the GPT-4 model was launched alongside its integration with Copilot X. Additionally, there was a further spike in November 2023, following the release of GPT-4 turbo. Overall, the trend diagram illustrates that the practice of self-declaring AI-generated code is on the rise as more advanced models are introduced. However, we see a sudden drop in the trend after April 2024, which may be attributed to the timing of our data collection from GitHub. Consequently, newer self-declared AI-generated code snippets may not have been fully indexed on GitHub yet or may not have gained sufficient visibility to be captured in our analysis.
@@ 7. Conclusions and Future Work
Our study also opens up several directions for future research. (1) We identify various reasons for self-declaring (or not) AI-generated code, which can be further investigated with a larger set of developers. Additionally, we provide guidelines for self-declaring AI-generated code by analyzing how developers report their generated code based on data from the mining study and practitioners’ survey. Future research can further refine and expand the guidelines using new datasets and validate their usefulness through empirical evaluation. (2) Our research shows that developers consider self-declaration comments beneficial for tracking AI-generated code for future review, debugging, and improvement. This opens up avenues for future research to investigate the impact of self-declaration comments on various code quality attributes, such as readability and maintainability. Such studies can contribute to a deeper understanding of the evolving landscape of software development, where AI-generated code is increasingly integrated into modern software systems. (3) It would be valuable to provide a code-commenting function in AI code generation tools to support the guidelines (see Section 5.3) for developers to declare AI-generated code.
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.

### v3

@@ 1. Introduction
Despite their usefulness and high potential (Wang et al., 2024; Liang et al., 2024), previous studies have shown that there are many quality (Yetiştiren et al., 2023; Liu et al., 2024b), correctness (Tambon et al., 2025; Vatai et al., 2024; Liu et al., 2024c), security (Pearce et al., 2022; Majdinasab et al., 2024; Fu et al., 2025) and ethical  (Stalnaker et al., 2024) challenges associated with AI-generated code (i.e., code that is primarily generated with AI coding assistance tools). These challenges stress the need to distinguish AI-generated code from other human-written code (Xu et al., 2025; Idialu et al., 2024) for code review and quality assurance purposes (Xu et al., 2025). To this end, we seek to study whether developers self-declare AI-generated code, explore the characteristics of self-declaration comments, and explore how and why developers self-declare (or not) AI-generated code. We derive our definition of self-declared AI-generated code from the well-established definition of self-admitted technical debt (Potdar and Shihab, 2014): a self-declared AI-generated code is code that has been explicitly acknowledged in the code comments by the developer as automatically generated by AI tools.
@@ 3. Research Methodology
Our study employs a mixed-methods approach comprising two main phases: (1) an analysis of self-declaration comments from AI-generated code collected from open-sourced projects, and (2) a survey with practitioners regarding the practices of self-declaring AI-generated code. We first collected data about AI-generated code from real-world projects by mining GitHub repositories to answer our research question of how (RQ1) developers self-declare AI-generated code. We followed that with a practitioner survey to validate the results from the mining study (RQ1) while addressing the research question of why (RQ2) developers self-declare AI-generated code. This methodological triangulation strengthens the validity of our findings by complementing the mining study results with developers’ perspectives from the practitioners’ survey. Figure 2 shows an overview of our research methodology.
@@ 1 Tool Names Only ChatGPT, GitHub Copilot, Code Llama, Amazon CodeWhisperer
In the pilot search phase, we collected a significant number of self-declared AI-generated code using search terms that included the names of the AI code generation tools and phrases specific to code generation. Therefore, out of the total 23 search terms finally selected for our data collection, 20 were based on Strategy 3 (see Table 1). The list included only 2 “thanking” terms from Strategy 4 because we also found relevant results using these two terms during the pilot phase. We only included one general search term without the AI tool name (Strategy 2) because we found limited results for such search terms. We could not find any results using the terms from Strategy 1, and we did not include any search terms from this strategy. Based on the results of the pilot search, 22 search terms included the names of the AI code generation tools. 17 out of the 22 search terms included names of ChatGPT and GitHub Copilot, and only 5 search terms included names of other tools. The list of search terms is provided in Table 2, which also shows the number of AI-generated code snippets included in our dataset.
@@ 3.2.2. Search and Filtration Process
We initially curated slightly over 1,200 code files from public GitHub repositories. After applying our inclusion and exclusion criteria (see Table 3), we narrowed down the selection to 899 files that fulfilled our criteria. We retained 613 code files from 586 unique repositories using manual sub-sampling to ensure a diverse dataset for data analysis. We included a few representative files from repositories containing multiple files with AI-generated code. This was necessary as several repositories contained many files with highly similar self-declaration comments. For instance, the repository KV6002-NUMUNCHIES alone contained 65 files with similar self-declaration comments with small variations. On the contrary, for the analysis of the distribution of AI-generated code in RQ1.3, we recorded the total number of files containing self-declared AI-generated code snippets in the “Total files with SDC” column (code-details.xlsx) in our dataset (Kashif et al., 2025).
@@ 3.2.4. Data Analysis
Initial Coding. We performed the initial coding by identifying and labeling essential ideas and concepts from self-declaration comments. The first author carefully read each self-declaration comment to extract codes. We iteratively refined the code as we examined more data. The other co-authors continuously reviewed the codes and shared their feedback to improve the reliability of the coding process. The first author shared the initial coding results with the co-authors through a collaborative document. The refinement process occurred in two rounds. First, the co-authors reviewed the document and provided written comments. Then, any remaining disagreements were discussed and resolved during live meetings. This approach aligns with established qualitative research practices, where a single primary coder conducts the initial coding while co-authors engage in iterative review and discussion to enhance the reliability of the coding process (McDonald et al., 2019). The analysis was done using MAXQDA [^2] https://www.maxqda.com/ to organize and manage the coding results.
@@ 3.3.3. Data Analysis
Analyzing Responses. We analyzed all responses to the survey questionnaire in Table 5 using a combination of quantitative and qualitative analysis methods. For the first nine closed-ended questions, we employed descriptive statistics to ascertain the frequency of various responses, which provides an overview of the participants’ demographics and their usage of AI-generated code. For the remaining three open-ended questions (SQ10, SQ11, SQ12), we employed thematic analysis and utilized the inductive approach for analysis (Braun and Clarke, 2006). Specifically, we first performed initial coding and then constantly compared the codes to develop high-level themes. The first author performed the initial coding by examining each valid response to extract and summarize the core ideas articulated in participants’ responses regarding their reasons to self-declare AI-generated code. The initial coding results were shared with the co-authors via a collaborative document. The refinement process took place in two phases. In the first phase, the co-authors reviewed the document and shared their written comments. Afterward, any disagreements were addressed during live meetings. This approach aligns with established qualitative research practices, where a single primary coder conducts the initial coding while co-authors engage in iterative review and discussion to enhance the reliability of the coding process (McDonald et al., 2019).
@@ 3.4. Mining Study Overview
Commit history. Figure 6 shows the commit history of the repositories that contain self-declared AI-generated code. We categorized repositories into six commit count ranges: from 1 – 50 to 5000. In total, we collected data from 586 repositories. 194 (33.1%) repositories are in the 100 - 500 commits range, indicating reasonable activity in these projects. Moreover, 137 repositories (23.4%) contain commit counts ranging from 1 to 50, showing certain inactivity in these projects. The projects containing more than 500 commits account for 136 (23.2%) repositories.
@@ 4.1. RQ1: How do developers self-declare AI-generated code in their projects?
1) Simple self-declaration. We identified 214 (34.9%) self-declaration comments as simple self-declaration comments. These comments self-declare AI-generated code in a few words, simply acknowledging the code is AI-generated. In most cases (over 96%), the comment will also include the name of the code generation tool used. The example (Comment 1) in Table 7 shows a developer mentioning that Copilot was used to generate the code. Only 3.3% of these comments just self-declared the AI-generated code without mentioning the tool name. For example, we can see in Comment 3 in Table 7 that a developer simply self-declared a function as AI-generated without mentioning any further details. Some self-declaration comments (4%) also acknowledged the simultaneous use of multiple AI code generation tools. In Comment 2 in Table 7, a developer mentioned GitHub Copilot and ChatGPT simultaneously for code generation.
@@ 7. Written by [developer name], 4/9/2024 with the assistance of GitHub Copilot (comment)
Trend of self-declaring AI-generated code among developers. We visualized the dates when self-declaration comments were added to the files, covering the period from January 2021 to December 2024. We observed an increasing trend in self-declaring AI-generated code among developers, as shown in Figure 14. The trend exhibits continuous growth, punctuated by sudden spikes. A noticeable increase is observed starting in November 2022, coinciding with the introduction of the GPT-3.5 model. Similarly, we observed another significant rise in March 2023, when the GPT-4 model was launched alongside its integration with Copilot X. Additionally, there was a further spike in November 2023, following the release of GPT-4 turbo. Overall, the trend diagram illustrates that the practice of self-declaring AI-generated code is on the rise as more advanced models are introduced. However, we observed a sudden drop in the trend after April 2024. This decline may be due to a few factors: (1) the timing of our data collection from GitHub, as newer self-declared AI-generated code snippets might not have been fully indexed or gained enough visibility to be included in our analysis; and (2) the increasing popularity of other AI code generation tools, including autonomous coding agents, which were not part of our search terms.
@@ 7. Conclusions and Future Work
Our study also opens up several directions for future research. (1) We identify various reasons for self-declaring (or not) AI-generated code, which can be further investigated with a larger set of developers. Additionally, we provide guidelines for self-declaring AI-generated code by analyzing how developers report their generated code based on data from the mining study and practitioners’ survey. Future research can further refine and expand the guidelines using new datasets and validate their usefulness through empirical evaluation. (2) Our research shows that developers consider self-declaration comments beneficial for tracking AI-generated code for future review, debugging, and improvement. This opens up avenues for future research to investigate the impact of self-declaration comments on various code quality attributes, such as readability and maintainability. Such studies can contribute to a deeper understanding of the evolving landscape of software development, where AI-generated code is increasingly integrated into modern software systems. (3) It would be valuable to provide a code-commenting function in AI code generation tools to support the guidelines (see Section 5.3) for developers to declare AI-generated code. (4) Our study focuses on the self-declaration of code snippets generated by AI code generation tools. With the growing popularity of autonomous coding agents capable of generating entire code files, future research can investigate the self-declaration of AI-generated code in the context of these coding agents.
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.